Emergency Condition
Display Panel 1 2 3 4
A 18 26 33 14
18 26 36 14
B 15 22 30 12
11 18 31 5
C 22 28 35 10
23 32 38 15
Two-way ANOVA: Time versus Panel, Condition
Source DF SS MS F P
Panel 2 144.542 144.542 25.32 .0000
Condition 3 1,647.12 549.042 96.18 .0000
Interaction 6 24.25 4.042 .71 .6498
Error 12 68.50 5.708
Total 23 2,028.96
Tabulated statistics: Panel, Condition
Rows: Panel Columns: Condition
1 2 3 4 All
A 15.00 26.00 34.50 13.00 22.13
B 13.00 20.00 30.50 7.50 17.75
C 22.50 30.00 36.50 16.00 26.25
All 17.67 25.33 33.83 12.17 22.04

  

35 30 25 20 15 10 3 Panel A Panel B Panel C

(e) Make pairwise comparisons of emergency conditions 1, 2, 3, and 4 by using Tukey simultaneous 95 percent confidence intervals. (Round your answers to 2 decimal places. Negative amounts should be indicated by a minus sign.)

u1u2: [ ,  ]
u1u3: [ ,  ]
u1u4: [ ,  ]
u2u3: [ ,  ]
u2u4: [ ,  ]
u3u4: [ ,  ]

(f) Which display panel minimizes the time required to stabilize an emergency condition? Does your answer depend on the emergency condition? Why?

(Click to select)Panel APanel BPanel C minimizes the time required to stabilize an emergency condition.
(Click to select)YesNo, there is (Click to select)someno interaction.

(g) Calculate a 95 percent (individual) confidence interval for the mean time required to stabilize emergency condition 4 using display panel B. (Round your answers to 2 decimal places.)

  Confidence interval         [ ,  ]
Answer #1

Answer:

MINITAB used.

The ANOVA results are different for this given data.

Analysis of Variance

Source

DF

Adj SS

Adj MS

F-Value

P-Value

Panel

2

228.58

114.292

19.32

0.000

condition

3

1651.00

550.333

93.01

0.000

Panel*condition

6

32.75

5.458

0.92

0.512

Error

12

71.00

5.917

Total

23

1983.33

(e) Make pairwise comparisons of emergency conditions 1, 2, 3, and 4 by using Tukey simultaneous 95 percent confidence intervals. (Round your answers to 2 decimal places. Negative amounts should be indicated by a minus sign.)

Tukey Simultaneous Tests for Differences of Means

Difference
of condition
Levels

Difference
of Means

SE of
Difference

Simultaneous
95% CI

T-Value

Adjusted
P-Value

2 - 1

7.50

1.40

(3.33, 11.67)

5.34

0.001

3 - 1

16.00

1.40

(11.83, 20.17)

11.39

0.000

4 - 1

-6.17

1.40

(-10.34, -2.00)

-4.39

0.004

3 - 2

8.50

1.40

(4.33, 12.67)

6.05

0.000

4 - 2

-13.67

1.40

(-17.84, -9.50)

-9.73

0.000

4 - 3

-22.17

1.40

(-26.34, -18.00)

-15.78

0.000

Individual confidence level = 98.83%

u1 u2:

[-11.67 ,-3.33  ]

u1 u3:

[-20.17,-11.83  ]

u1 u4:

[2.00 , 10.34 ]

u2 u3:

[ -12.67, -4.33 ]

u2 u4:

[9.50 ,17.84  ]

u3 u4:

[ 18.00, 26.34 ]

(f) Which display panel minimizes the time required to stabilize an emergency condition? Does your answer depend on the emergency condition? Why?

(Click to select)Panel BPanel minimizes the time required to stabilize an emergency condition.
(Click to select)No, there is (Click to select)No interaction.

(g) Calculate a 95 percent (individual) confidence interval for the mean time required to stabilize emergency condition 4 using display panel B. (Round your answers to 2 decimal places.)

  Confidence interval         [ -7.19, 24.19 ]

Mean=8.5 and se= 7.2

T table value at 0.05 level=2.179

Lower limit= 8.5-2.179*7.2 = -7.1888

upper limit= 8.5+2.179*7.2 = 24.1888

Means

Term

Fitted
Mean

SE Mean

Panel

A

23.125

0.860

B

18.000

0.860

C

25.375

0.860

condition

1

17.833

0.993

2

25.333

0.993

3

33.833

0.993

4

11.667

0.993

Panel*condition

A 1

18.00

1.72

A 2

26.00

1.72

A 3

34.50

1.72

A 4

14.00

1.72

B 1

13.00

1.72

B 2

20.00

1.72

B 3

30.50

1.72

B 4

8.50

1.72

C 1

22.50

1.72

C 2

30.00

1.72

C 3

36.50

1.72

C 4

12.50

1.72